### v1

@@ 3 Nyström approximation in infinite dimensions
This section presents an infinite-dimensional extension of the randomized Nyström approximation. We begin by briefly introducing the concepts of quasimatrices, Hilbert–Schmidt operators, and Gaussian processes, which will be useful to generalize the bounds of Section 2 to operators between function spaces. While we consider operators between square-integrable functions as the most important application, the results of this section extend naturally to other separable Hilbert spaces following the theory in [22].
@@ 3.1 Quasimatrices
For a bounded domain $D\subset\mathbb{R}^{d}$, $d\geq 1$, we consider the Hilbert space $L^{2}(D)$ of square-integrable functions. Quasimatrices are a convenient way to represent and work with collections of functions or more, generally, elements of infinite-dimensional vector spaces; see, e.g., [37]. In particular, a function $Y:\mathbb{R}^{m}\to L^{2}(D)$ is expressed as the quasimatrix
$Y=\begin{bmatrix}y_{1}&\cdots&y_{m}\end{bmatrix},\quad y_{i}\in L^{2}(D).$
@@ 3.2 Non-negative self-adjoint trace-class operators
We consider a non-negative self-adjoint trace-class operator $\mathcal{A}:L^{2}(D)\to L^{2}(D)$, i.e., it holds that $\langle\mathcal{A}f,f\rangle\geq 0$ and $\langle\mathcal{A}f,g\rangle=\langle f,\mathcal{A}g\rangle$ for every $f,g\in L^{2}(D)$, and the trace norm [22, Def. 4.5.1] is finite:
@@
for any orthonormal basis $\{e_{j}\}_{j}$ of $L^{2}(D)$. Non-negative self-adjoint trace-class operators are Hilbert–Schmidt operators [22, Thm. 4.5.2], and therefore admit an eigenvalue decomposition of the form [22, Thm. 4.3.1]:
$\mathcal{A}=\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}>0\end{subarray}}^{\infty}\sigma_{j}\langle u_{j},\cdot\rangle u_{j},$ (18)
where $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$ are the eigenvalues of $\mathcal{A}$, and $\{u_{j}\}_{j}$ are orthonormal eigenfunctions. The eigenvalues allow us to express the trace, Hilbert–Schmidt, and operator norms of $\mathcal{A}$ as
@@
which are infinite-dimensional analogs of the nuclear, Frobenius, and spectral norms discussed in Section 2. Furthermore, we introduce the operator $U:\ell^{2}\to L^{2}(D)$ defined by $Uf=\sum_{i=1}^{\infty}f_{i}u_{i}$ for any $f$ in $\ell^{2}$, the space of square-summable sequences (indexed by positive integers). Then, for a given rank $k\geq 1$, the quasimatrix $U_{1}:\mathbb{R}^{k}\to L^{2}(D)$ contains the first $k$ eigenfunctions of $\mathcal{A}$, and the quasimatrix $U_{2}:\ell^{2}\to L^{2}(D)$ contains the remaining eigenfunctions. Finally, we introduce the diagonal matrix and quasimatrices $\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k})$ and $\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\sigma_{k+2},\ldots)$, respectively, which contain the eigenvalues of $\mathcal{A}$ in descending order.
@@ 3.3 Gaussian processes
Let $K:D\times D\mapsto\mathbb{R}$ be a continuous symmetric positive semi-definite kernel, with the following eigenvalue decomposition [22, Thm. 4.6.5]:
@@
$\mathcal{K}[f](x)=\int_{D}K(x,y)f(y)\,\mathrm{d}y,\quad f\in L^{2}(D),\,x\in D,$
and $\{\psi_{j}\}_{j}$ are the corresponding orthonormal eigenfunctions of $\mathcal{K}$ in $L^{2}(D)$. In the following, we assume that $\mathcal{K}$ is trace-class, i.e., $\sum_{i=1}^{\infty}\lambda_{i}<\infty$. A Gaussian process (GP) $\omega\sim\mathcal{GP}(0,K)$ is a stochastic process such that for any finite collection of points $x_{1},\ldots,x_{m}\in D$, the vector $\begin{bmatrix}\omega(x_{1})&\cdots&\omega(x_{m})\end{bmatrix}^{\top}$ follows a multivariate normal distribution with mean $\bm{0}\in\mathbb{R}^{m}$ and covariance matrix $\bm{K}=(K(x_{i},x_{j}))_{1\leq i,j\leq m}$. The Karhunen–Loève expansion [1, Ch.3.2] of $\omega$ is given by
$\omega\sim\sum_{k=1}^{\infty}\lambda_{k}^{1/2}\zeta_{k}\psi_{k},$ (19)
where $\zeta_{1},\zeta_{2},\ldots\sim\mathcal{N}(0,1)$ are mutually independent. With probability one, a realization of $\omega$ is in $L^{2}(D)$ [22, Thm. 7.2.5]. Recalling that $\{u_{j}\}_{j}$ denote the eigenfunctions of $\mathcal{A}$, see (18), we define the function $\widetilde{\bm{K}}:\mathbb{N}^{*}\times\mathbb{N}^{*}\to\mathbb{R}$ elementwise as
$\widetilde{\bm{K}}(i,j)=\langle u_{i},\mathcal{K}u_{j}\rangle=\sum_{k=1}^{\infty}\lambda_{k}\langle\psi_{k},u_{i}\rangle\langle\psi_{k},u_{j}\rangle,\quad i,j\in\mathbb{N}^{*},$
which is bounded by $\|\mathcal{K}\|_{\mathrm{Tr}}$ using the Cauchy–Schwarz inequality. We also define the following restrictions of $\widetilde{\bm{K}}$ to different sets of indices:
@@
in analogy to Eq. 3.
@@ 3.4 Infinite-dimensional extension of the Nyström approximation
$\widehat{\mathcal{A}}:=\mathcal{A}\Omega(\Omega^{*}\mathcal{A}\Omega)^{\dagger}(\mathcal{A}\Omega)^{*}.$ (20)
Assuming that the realization of $\omega_{i}$ is in $L^{2}(D)$, which holds with probability $1$, the Nyström approximation is an operator $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ of rank at most $k+p$ with the explicit representation
$\widehat{\mathcal{A}}[f]=\sum_{i,j=1}^{k+p}\mathcal{A}\omega_{i}\big{[}(\Omega^{*}\mathcal{A}\Omega)^{\dagger}\big{]}_{ij}\langle\omega_{j},\mathcal{A}f\rangle.$
As in the finite-dimensional case, the error bounds for the infinite-dimensional analog of the Nyström approximation depend on the prior information of eigenvectors contained in $\mathcal{K}$, which is measured by the following two quantities:
$\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-1}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},$ (21)
where $\xi\in\{\mathrm{HS},\mathrm{Tr},\mathrm{op}\}$. These quantities are the infinite-dimensional analogs of Eq. 4.
For each $1\leq j\leq k+p$, we consider the stochastic process $\bm{\omega}_{j}=\{\langle u_{i},\omega_{j}\rangle,\,i\in\mathbb{N}^{*}\}$ whose trajectories are in $\ell^{2}$, and denote by $\bm{\Omega}=U^{*}\Omega=\begin{bmatrix}\bm{\omega}_{1}&\cdots&\bm{\omega}_{k+p}\end{bmatrix}$ the random quasimatrix whose columns are i.i.d. from $\mathcal{GP}(0,\widetilde{\bm{K}})$. Then, we introduce the random $k\times(k+p)$ matrix $\bm{\Omega}_{1}$ as
@@ 3.4.1 Structural bound
Let $\bm{B},\bm{C}:\ell^{2}\to\ell^{2}$ be non-negative self-adjoint trace-class operators such that $\bm{B}-\bm{C}$ is non-negative. For $n\in\mathbb{N}^{*}$, consider the restriction $\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}$ of $\bm{B}$ to its first $n$ rows and columns. Then, for $\xi\in\{\mathrm{Tr},\mathrm{HS},\mathrm{op}\}$, we have
$\lim_{n\rightarrow\infty}\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}$ $=\|\bm{B}\|_{\xi},$ (22a)
$\|\bm{C}\|_{\xi}\leq\|\bm{B}\|_{\xi}$ $\leq\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}+\|\bm{B}_{\llbracket n+1,\infty)\times\llbracket n+1,\infty)}\|_{\xi}.$ (22b)
For $\xi\in\{\mathrm{HS},\mathrm{Tr}\}$, the property (22a) follows from the absolute convergence of the involved series. For $\xi=\mathrm{op}$, the triangle inequality yields
$|\|\bm{B}\|_{\mathrm{op}}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{op}}|\leq\sqrt{\|\bm{B}\|_{\mathrm{HS}}^{2}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{HS}}^{2}}\rightarrow 0,\quad\text{as }n\to\infty.$
Eq. Eq. 22b is an infinite-dimensional analog of [25, Thm. 2.1] and [32, Lem. 3.1], which can be proven using Eq. 22a. We can now proceed with an infinite-dimensional analog of the finite-dimensional structural bound Eq. 7.
@@
where the third equality follows from the unitary invariance of the norms and the fourth equality is due to the relation $U^{*}\mathcal{A}U=\bm{\Sigma}$. As in the finite-dimensional case, the rest of the proof follows the argument of the proof of [32, Lem. 3.13] for the operator monotone function $f:x\mapsto x$ and $q=1$, now using the fact that the inequalities in Eq. 22b are infinite-dimensional analogs of [32, Lems. A.1 and 3.1] with $f:x\mapsto x$ and $\|\cdot\|\coloneqq\|\cdot\|_{\xi}$.

### v2

@@ 3 Nyström approximation in infinite dimensions
This section presents an infinite-dimensional extension of the randomized Nyström approximation. We begin by briefly introducing the concepts of quasimatrices, Hilbert-Schmidt operators, and Gaussian processes, which will be useful to generalize the bounds of Section 2 to operators between function spaces. While we consider operators between square-integrable functions as the most important application, the results of this section extend naturally to other separable Hilbert spaces following the theory in [24].
@@ 3.1 Quasimatrices
For a bounded domain $D\subset\mathbb{R}^{d}$, $d\geq 1$, we consider the Hilbert space $L^{2}(D)$ of real-valued square-integrable functions. Quasimatrices are a convenient way to represent and work with collections of functions or more, generally, elements of infinite-dimensional vector spaces; see, e.g., [38]. In particular, a linear operator $Y:\mathbb{R}^{m}\to L^{2}(D)$ can be expressed as the quasimatrix
(20) $Y=\begin{bmatrix}y_{1}&\cdots&y_{m}\end{bmatrix},\quad y_{i}\in L^{2}(D){\color[rgb]{0,0,0},}$
@@
$Y^{*}f=\begin{bmatrix}\langle y_{1},f\rangle\\ \vdots\\ \langle y_{m},f\rangle\end{bmatrix}\in\mathbb{R}^{m}.$
If $Z:\mathbb{R}^{\ell}\to L^{2}(D)$ is another quasimatrix, then $Y^{*}Z$ yields a linear operator $\mathbb{R}^{\ell}\to\mathbb{R}^{m}$, which with respect to the the canonical basis has the representing matrix:
@@
Furthermore, given a linear operator $\mathcal{B}:L^{2}(D)\to L^{2}(D)$, we write $\mathcal{B}Y$ to denote the quasimatrix whose $i^{\text{th}}$ column is $\mathcal{B}y_{i}$.
Our discussion so far has focused on the case where quasimatrices have a finite number of columns. In general, the definition of a quasimatrix cannot be extended to cases where the quasimatrix has a countably infinite number of columns. However, if the columns form an orthonormal system, it is possible to extend the definition. Consider an orthonormal system $\{u_{i}\}_{i\in\mathbb{N}}\in L^{2}(D)$ and let $\ell^{2}$ denote the set of square-summable sequences. The quasimatrix $U=\begin{bmatrix}u_{1}&u_{2}&\cdots\end{bmatrix}$ is defined to be the bounded linear operator $\ell^{2}\to L^{2}(D)$ satisfying $U\bm{e}_{i}=u_{i}$, where $\bm{e}_{i}$ is the $i^{\text{th}}$ canonical vector in $\ell^{2}$. Furthermore, for a Hilbert-Schmidt operator $\mathcal{B}:L^{2}(D)\to L^{2}(D)$ we define $\mathcal{B}U:\ell^{2}\to L^{2}(D)$ as the linear operator satisfying $\mathcal{B}U\bm{e}_{i}=\mathcal{B}u_{i}$. One can verify that $\mathcal{B}U$ is also a Hilbert-Schmidt operator [24, Def. 4.4.4].
@@ 3.2 Non-negative self-adjoint trace-class operators
We consider a non-negative self-adjoint trace-class operator $\mathcal{A}:L^{2}(D)\to L^{2}(D)$, i.e., it holds that $\langle\mathcal{A}f,f\rangle\geq 0$ and $\langle\mathcal{A}f,g\rangle=\langle f,\mathcal{A}g\rangle$ for every $f,g\in L^{2}(D)$, and the trace norm [24, Def. 4.5.1] is finite:
@@
for any orthonormal basis $\{e_{j}\}_{j}$ of $L^{2}(D)$. Non-negative self-adjoint trace-class operators are Hilbert-Schmidt operators [24, Thm. 4.5.2], and therefore admit an eigenvalue decomposition of the form [24, Thm. 4.3.1]:
(21) $\mathcal{A}=\sum_{\begin{subarray}{c}j=1\\ \sigma_{j}>0\end{subarray}}^{\infty}\sigma_{j}\langle u_{j},\cdot\rangle u_{j},$
where $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$ are the eigenvalues of $\mathcal{A}$, and $\{u_{j}\}_{j}$ are orthonormal eigenfunctions forming a complete orthonormal system for the closure of the image of $\mathcal{A}$ [24, Theorem 4.2.4]. The series (21) is converging in the operator norm; see [24, p.105]. The eigenvalues allow us to express the trace, Hilbert-Schmidt, and operator norms of $\mathcal{A}$ as
@@
which are infinite-dimensional analogs of the nuclear, Frobenius, and spectral norms discussed in Section 2. Furthermore, we introduce the operator $U:\ell^{2}\to L^{2}(D)$ defined by $Uf=\sum_{i=1}^{\infty}f_{i}u_{i}$ for any $f$ in $\ell^{2}$, the space of square-summable sequences (indexed by positive integers). Then, for a given rank $k\geq 1$, the quasimatrix $U_{1}:\mathbb{R}^{k}\to L^{2}(D)$ contains the first $k$ eigenfunctions of $\mathcal{A}$, and the quasimatrix $U_{2}:\ell^{2}\to L^{2}(D)$ contains the remaining eigenfunctions. In addition, the adjoint of $U$ is the operator $U^{*}:L^{2}(D)\to\ell^{2}$ defined $(U^{*}g)_{i}=\langle u_{i},g\rangle$ for any $g\in L^{2}(D)$, where $(U^{*}g)_{i}$ is the $i^{\text{th}}$ entry in the sequence $U^{*}g$. $U_{1}^{*}$ and $U_{2}^{*}$ can be defined in an analogous way. Finally, we introduce the diagonal matrix and infinite diagonal matrix $\bm{\Sigma}_{1}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{k})$ and $\bm{\Sigma}_{2}=\operatorname{diag}(\sigma_{k+1},\sigma_{k+2},\ldots)$, respectively, which contain the eigenvalues of $\mathcal{A}$ in descending order.
@@ 3.3 Gaussian processes
Let $K:D\times D\mapsto\mathbb{R}$ be a continuous symmetric positive semi-definite kernel, with the following eigenvalue decomposition [24, Thm. 4.6.5]:
@@
$[\mathcal{K}f](x)=\int_{D}K(x,y)f(y)\,\mathrm{d}y,\quad f\in L^{2}(D),\,x\in D,$
and $\{\psi_{j}\}_{j}$ are the corresponding orthonormal eigenfunctions of $\mathcal{K}$ in $L^{2}(D)$. In the following, we assume that $\mathcal{K}$ is trace-class, i.e., $\sum_{i=1}^{\infty}\lambda_{i}<\infty$. A Gaussian process (GP) $\omega\sim\mathcal{GP}(0,K)$ is a stochastic process such that for any finite collection of points $x_{1},\ldots,x_{m}\in D$, the vector $\begin{bmatrix}\omega(x_{1})&\cdots&\omega(x_{m})\end{bmatrix}^{\top}$ follows a multivariate normal distribution with mean $\bm{0}\in\mathbb{R}^{m}$ and covariance matrix $\bm{K}=(K(x_{i},x_{j}))_{1\leq i,j\leq m}$. The Karhunen-Loève expansion [1, Ch.3.2] of $\omega$ is given by
(22) $\omega\sim\sum_{k=1}^{\infty}\lambda_{k}^{1/2}\zeta_{k}\psi_{k},$
where $\zeta_{1},\zeta_{2},\ldots\sim\mathcal{N}(0,1)$ are mutually independent. With probability one, a realization of $\omega$ is in $L^{2}(D)$ [24, Thm. 7.2.5]. Recalling that $\{u_{j}\}_{j}$ denote the eigenfunctions of $\mathcal{A}$, see (21), we define the linear operator $\widetilde{\bm{K}}:\ell^{2}\to\ell^{2}$, which under the canonical basis has the following infinite matrix representation
${\color[rgb]{0,0,0}\widetilde{\bm{K}}_{i,j}}=\langle u_{i},\mathcal{K}u_{j}\rangle=\sum_{k=1}^{\infty}\lambda_{k}\langle\psi_{k},u_{i}\rangle\langle\psi_{k},u_{j}\rangle,\quad i,j\in\mathbb{N}^{*}.$
Note that $\widetilde{\bm{K}}$ is also a trace-class operator. We also define the following restrictions of $\widetilde{\bm{K}}$ to different sets of indices:
@@
in analogy to (3). Similar to the discussion above, one can define $\mathcal{GP}(0,\widetilde{\bm{K}})$ as follows: a stochastic process $\bm{\omega}$ in $\ell^{2}$ is said to be distributed as $\mathcal{GP}(0,\widetilde{\bm{K}})$ if, for any finite collection of indices $\ell_{1},\ldots,\ell_{m}\in\mathbb{N}^{*}$, the vector $\begin{bmatrix}\bm{\omega}_{\ell_{1}}&\cdots&\bm{\omega}_{\ell_{m}}\end{bmatrix}$ follows a multivariate normal distribution with mean $\bm{0}\in\mathbb{R}^{m}$ and covariance matrix $\bm{K}=(\widetilde{\bm{K}}_{\ell_{i},\ell_{j}})_{1\leq i,j\leq m}$. [^1] By the Kolmogorov two-series theorem, $\bm{\omega}\in\ell^{2}$ almost surely.
@@ 3.4 Infinite-dimensional extension of the Nyström approximation
(23) $\widehat{\mathcal{A}}:=\mathcal{A}\Omega(\Omega^{*}\mathcal{A}\Omega)^{\dagger}(\mathcal{A}\Omega)^{*}.$
Assuming that the realization of $\omega_{i}$ is in $L^{2}(D)$, which holds with probability $1$, the Nyström approximation is an operator $\widehat{\mathcal{A}}:L^{2}(D)\to L^{2}(D)$ of rank at most $k+p$ whose action on a function $f\in L^{2}(D)$ has the explicit representation
${\color[rgb]{0,0,0}[\widehat{\mathcal{A}}f](x)=\sum_{i,j=1}^{k+p}[\mathcal{A}\omega_{i}](x)\big{[}(\Omega^{*}\mathcal{A}\Omega)^{\dagger}\big{]}_{ij}\langle\omega_{j},\mathcal{A}f\rangle,}$
where $\big{[}(\Omega^{*}\mathcal{A}\Omega)^{\dagger}\big{]}_{ij}$ is the $(i,j)$-entry of the matrix $(\Omega^{*}\mathcal{A}\Omega)^{\dagger}$. As in the finite-dimensional case, the error bounds for the infinite-dimensional analog of the Nyström approximation depend on the prior information of eigenvectors contained in $\mathcal{K}$, which is measured by the following two quantities:
(24) $\beta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{22.1}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2},\quad{\color[rgb]{0,0,0}\delta_{k}^{(\xi)}=\frac{\|\bm{\Sigma}_{2}^{1/2}\widetilde{\bm{K}}_{21}\widetilde{\bm{K}}_{11}^{-2}\widetilde{\bm{K}}_{21}^{*}\bm{\Sigma}_{2}^{1/2}\|_{\xi}}{\|\bm{\Sigma}_{2}\|_{\xi}}},$
[…]
For each $1\leq j\leq k+p$, we consider the stochastic process $\bm{\omega}_{j}=\{\langle u_{i},\omega_{j}\rangle,\,i\in\mathbb{N}^{*}\}$ whose trajectories are in $\ell^{2}$, and denote by $\bm{\Omega}=U^{*}\Omega=\begin{bmatrix}\bm{\omega}_{1}&\cdots&\bm{\omega}_{k+p}\end{bmatrix}$ the random infinite matrix whose columns are i.i.d. from $\mathcal{GP}(0,\widetilde{\bm{K}})$. Then, we introduce the random $k\times(k+p)$ matrix $\bm{\Omega}_{1}$ as
@@ 3.4.1 Structural bound
Let $\bm{B},\bm{C}:\ell^{2}\to\ell^{2}$ be non-negative self-adjoint trace-class operators such that $\bm{B}-\bm{C}$ is non-negative. For $n\in\mathbb{N}^{*}$, consider the restriction $\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}$ of $\bm{B}$ to its first $n$ rows and columns under the canonical basis of $\ell^{2}$. Then, for $\xi\in\{\mathrm{Tr},\mathrm{HS},\mathrm{op}\}$, we have
(25a) $\lim_{n\rightarrow\infty}\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}$ $=\|\bm{B}\|_{\xi},$
(25b) $\|\bm{C}\|_{\xi}\leq\|\bm{B}\|_{\xi}$ $\leq\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}+\|\bm{B}_{\llbracket n+1,\infty)\times\llbracket n+1,\infty)}\|_{\xi}.$
Proof 3.2.
For $\xi\in\{\mathrm{HS},\mathrm{Tr}\}$, the property (25a) follows from the absolute convergence of the involved series $\lim_{n\rightarrow\infty}\sum_{i,j=1}^{n}\bm{B}_{ij}^{2}=\sum_{i,j=1}^{\infty}\bm{B}_{ij}^{2}$ and $\lim_{n\rightarrow\infty}\sum_{i=1}^{n}\bm{B}_{ii}=\sum_{i=1}^{\infty}\bm{B}_{ii}$, where the entries $\bm{B}_{ij}$ are defined with respect to the canonical basis of $\ell^{2}$. For $\xi=\mathrm{op}$, let $\bm{P}_{n}$ be the orthogonal projection onto the first $n$ canonical basis vectors in $\ell^{2}$. Then, $\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{op}}=\|\bm{P}_{n}\bm{B}\bm{P}_{n}\|_{\mathrm{op}}$. Hence, the triangle inequality yields
$|\|\bm{B}\|_{\mathrm{op}}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{op}}|$ $\leq\|\bm{B}-\bm{P}_{n}\bm{B}\bm{P}_{n}\|_{\mathrm{op}}$
$\leq\|\bm{B}-\bm{P}_{n}\bm{B}\bm{P}_{n}\|_{\mathrm{HS}}$
$=\sqrt{\|\bm{B}\|_{\mathrm{HS}}^{2}-\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\mathrm{HS}}^{2}}\rightarrow 0,\quad\text{as }n\to\infty.$
To obtain (25b), first pick some $m\geq n$. Combining [27, Thm. 2.1] and [34, Lem. 3.1], we have
$\|\bm{C}_{\llbracket 1,m\rrbracket\times\llbracket 1,m\rrbracket}\|_{\xi}\leq\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}\leq\|\bm{B}_{\llbracket 1,n\rrbracket\times\llbracket 1,n\rrbracket}\|_{\xi}+\|\bm{B}_{\llbracket n+1,m)\times\llbracket n+1,m)}\|_{\xi}.$
Taking the limit as $m\to\infty$ and using (25a) yields the desired result.
We can now proceed with an infinite-dimensional analog of the finite-dimensional structural bound (5).
@@
where the third equality follows from the unitary invariance of the norms and the fourth equality is due to the relation $U^{*}\mathcal{A}U=\bm{\Sigma}$. Now note that $\|\bm{\Sigma}-\bm{\Sigma}\bm{\Omega}(\bm{\Omega}^{*}\bm{\Sigma}\bm{\Omega})^{\dagger}(\bm{\Sigma}\bm{\Omega})^{*}\|_{\xi}=\|\bm{\Sigma}^{1/2}(\bm{I}-\bm{P}_{\bm{\Sigma}^{1/2}\bm{\Omega}})\bm{\Sigma}^{1/2}\|_{\xi},$ where $\bm{I}$ is the identity operator and $\bm{P}_{\bm{\Sigma}^{1/2}\bm{\Omega}}$ is the orthogonal projection onto $\operatorname{range}(\bm{\Sigma}^{1/2}\bm{\Omega})$. Let
$\bm{Z}=\bm{\Sigma}\bm{\Omega}\bm{\Omega}_{1}^{\dagger}\bm{\Sigma}_{1}^{-1/2}=\begin{bmatrix}\bm{I}\\ \bm{F}\end{bmatrix},\quad\bm{F}=\bm{\Sigma}_{2}^{1/2}\bm{\Omega}_{2}\bm{\Omega}_{1}^{\dagger}\bm{\Sigma}_{1}^{-1/2},$
where we used that $\bm{\Omega}_{1}$ is full rank. We have $\operatorname{range}(\bm{Z})\subseteq\operatorname{range}(\bm{\Sigma}^{1/2}\bm{\Omega})$ and therefore $\bm{I}-\bm{P}_{\bm{Z}}\succeq\bm{I}-\bm{P}_{\bm{\Sigma}^{1/2}\bm{\Omega}}$, where $\bm{P}_{\bm{Z}}=\bm{Z}(\bm{Z}^{*}\bm{Z})^{\dagger}\bm{Z}^{*}$ is the orthogonal projection onto $\operatorname{range}(\bm{Z})$. Hence, by Lemma 3.1 $\|\bm{\Sigma}^{1/2}(\bm{I}-\bm{P}_{\bm{\Sigma}^{1/2}\bm{\Omega}})\bm{\Sigma}^{1/2}\|_{\xi}\leq\|\bm{\Sigma}^{1/2}(\bm{I}-\bm{P}_{\bm{Z}})\bm{\Sigma}^{1/2}\|_{\xi}$. As in the finite-dimensional case, the rest of the proof follows the argument of the proof of [34, Lem. 3.13] for the operator monotone function $f:x\mapsto x$ and $q=1$, now using the $2\times 2$ block-structure of $\bm{I}-\bm{P}_{\bm{Z}}$:
$\bm{I}-\bm{P}_{\bm{Z}}=\begin{bmatrix}\bm{I}-(\bm{I}+\bm{F}^{*}\bm{F})^{-1}&-(\bm{I}+\bm{F}^{*}\bm{F})^{-1}\bm{F}^{*}\\ -\bm{F}(\bm{I}+\bm{F}^{*}\bm{F})^{-1}&\bm{I}-\bm{F}(\bm{I}+\bm{F}^{*}\bm{F})^{-1}\bm{F}^{*}\end{bmatrix}$
and the fact that the inequalities in (25b) are infinite-dimensional analogs of [34, Lems. A.1 and 3.1] with $f:x\mapsto x$ and $\|\cdot\|\coloneqq\|\cdot\|_{\xi}$.